OpenStudy (anonymous):

Katie invested $1200 at 8.5% interest. She earned $612 interest at the end of y years. What is the value of y? A. 18 B. 12 C. 6 D. 3

OpenStudy (anonymous):

http://openstudy.com/users/jaynator495

OpenStudy (anonymous):

@Jaynator495

OpenStudy (anonymous):

5 MEDALS

OpenStudy (anonymous):

SO U WANT TO TAKE THE CHALANGE

OpenStudy (anonymous):

ans is C INTEREST = PRT/100 612= (1200*8.5*T) / 100 THEREFORE T = 6 which is the number of years

OpenStudy (anonymous):

THANX CAN U HELP ME AGAIN